2021-03-22

Observers: Timo Kravtsov , Claudia Gutiérrez, Steve Williams

Support: Takashi Nagao

TAT: Tassilo Schweyer, Jesper Sollerman, Sheng Yang

EFOSC2 Calibrations: Darks, bias frames, Dome Flats, Arcs

Weather:

Start: humidity 37%, wind 2.9 m/s, seeing 1.2

Middle: humidity 34%, wind 3.2 m/s, seeing 1.2

End: humidity 38%, wind 2.8 m/s, seeing 1.3

Comments:

During the day, the engineers were working in the issue reported last night, because of this we only got the DARKS. The rest of the calibrations were taken in the morning.

23:48: The image analysis was not possible, but some corrections were performed (focus).